Circular No. 6803

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J0053–724 AND AX J0051–722 R. Corbet, Goddard Space Flight Center (GSFC) and Universities

Space Research Association (USRA); F. E. Marshall, GSFC; J. C. Lochner, GSFC and USRA; and M. Ozaki and Y. Ueda, Institute of Space and Astronautical Science, report: “Further analysis of the ASCA observation of 1WGA J0053.8–7226 = XTE J0053–724 (IAUC 6788, 6789) reveals that there are two transient pulsing sources in this image. 1WGA J0053.8– 7226, whose location was given on IAUC 6788, in fact has a period of 46.63 ± 0.04 s, rather than ∼ 92 s as initially reported. A second source, AX J0051–722, located at α = 0^{h}51^{m}04^{s}, δ = −72^{o}13′44′′ (equinox 2000.0; ± 1′.2, 90-percent confidence) with a 2–10-keV flux of 1.05 × 10^{−11} erg s^{−1} cm^{−2}, exhibits the 91.12-s (± 0.05 s) pulsations first detected with RXTE. Initial misidentification arose due to the approximate 2 to 1 ratio of these periods. Additional analysis of the first RXTE observation (IAUC 6777) shows, in addition to the 46.6-s and 91.1-s pulsations, an indication of additional modulation at a period of 74.8 ± 0.4 s from a possible third pulsar within the 2^{o} FWZI field-of-view centered on SMC X-3. However, no corresponding ASCA source can be identified. A ROSAT observation on 1997 Dec. 11 detected both AX J0051–722 and 1WGA J0053.8–7226 at α = 0^{h}50^{m}55^{s}.8, δ = −72^{o}13′38′′, and at α = 0^{h}53^{m}55^{s}.0, δ = −72^{o}26′47′′ (± 10′′), respectively.”

GRB 971227 S. A. Ilovaisky and C. Chevalier, Observatoire de Haute-Provence

(OHP), write: “R-band CCD images of the SAX error box reported on IAUC 6797 were obtained with the 1.2-m OHP telescope on Jan. 3.14 UT. Analysis of the resulting median frame of three 30-min exposures shows no object at the position reported by Castro-Tirado et al. (IAUC 6800) down to a limiting R_{c} magnitude slightly fainter than 22. A group of faint galaxies with R_{c} ∼ 21.5 appears a few arcsec northwest of that position.”

SUPERNOVAE S. Moretti and S. Tomaselli, Forlí, Italy, report unfiltered CCD mag-

nitudes (GSC reference stars): SN 1997dn in NGC 3451, 1997 Dec. 29.94 UT, 16.7; SN 1997dq in NGC 3810, Dec. 29.9, 16.4; SN 1997ef in UGC 4107, Dec. 29.85, 16.8.

1998 January 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green